### v1

@@ Appendix B Evaluation Protocol
For testing, some datasets provide bounding boxes (e.g., SSP-3D), while others do not (e.g., AGORA). In the latter case, we use the YOLOv8 [39] detector (trained on COCO).
@@ Appendix C Training Details
The skeleton-based datasets do not provide their correspondence to our SMPL-compatible canonical space. Hence, we perform an approximate initialization and then treat the canonical position of each of several hundred joints (across the skeleton types) as trainable variables during the main training. The corresponding left and right joints are forced to be symmetrically placed in canonical space. To avoid divergence, we impose a constraint that the canonical positions cannot move away more than 70 mm from their initial positions.
@@
Appendix D Uncertainty Estimation
@@
[…]
where is the ?stop gradient? operation that blocks the gradient flow during backpropagation. Intuitively, for , multiplication by ?cancels out? the division of by , resulting in the same gradients for as in the case of vanilla Euclidean loss (removing the above-mentioned downweighting of challenging examples), without impacting the gradients for (due to ).
@@
Appendix E Efficient Body Model Fitting
@@
To increase efficiency further, we may use only a subset of the vertices of the body model during fitting. In Fig. 6, we can see that, e.g., by fitting to only 1024 vertices, we increase the throughput by a factor of 3, while mIoU only decreases from 0.845 to 0.843. The timing was measured on an NVIDIA RTX 3090 GPU with a batch size of 2048.
In the model transfer experiment described in the main paper, where we fit SMPL-X parameters to clean SMPL mesh vertices with our fitting algorithm, we do not regularize the betas (as the input is a clean mesh without noise), use one iteration and 4096 vertices for fitting (out of the 10475).
@@
We will make all code publicly available, including the fitting algorithm, to ensure its precise reproducibility.
Appendix F Neural Field Architecture and Details
@@
Appendix G Additional 3DPW Protocol
@@
Appendix H Further Qualitative Results

### v2

@@ Appendix B Evaluation Protocol
For testing, some datasets provide bounding boxes (e.g., SSP-3D), while others do not (e.g., AGORA). In the latter case, we use the YOLOv8 [42] detector (trained on COCO). In the case of AGORA, we evaluate in the fine-tuned setting (NLF fine-tuned just on AGORA), and here we also fine-tune the YOLOv8 detector on AGORA-train.
Evaluation metrics
In our evaluations, we apply the conventionally used metrics for each individual benchmark. MPJPE is the average Euclidean joint error after alignment at the pelvis. (All distance-based errors are given as millimeters in this work.) Following prior works, e.g., [101], on 3DPW and EMDB the reference point is the midpoint between the two hip joints. P-MPJPE is the Euclidean error after Procrustes alignment, i.e., least-squares optimal rigid and uniform scaling alignment. When 14 joints are indicated for evaluation, these are obtained with a Human3.6M-like regressor from [51], which is commonly used in the literature.
MVE is analogous to MPJPE but for mesh vertices instead of joints, similarly P-MVE is the Procrustes-aligned version of it.
MPJAE and P-MPJAE are the mean per joint angle errors without and with Procrustes alignment, in degrees. It is the geodesic distance on the rotation manifold, comparing the global body part orientations between ground truth and prediction.
On SSP-3D, PVE-T-SC evaluates the body shape, by computing the average Euclidean error for the predicted and the ground truth vertices in the default T-pose, after the prediction is optimally scale-aligned to the ground truth. The mIoU is the mean intersection over union metric comparing the binary masks obtained through projecting the ground truth mesh and the predicted mesh on the image plane.
On AGORA, NMVE is a normalized version of MVE that takes into account the detection performance as well. It is the product of MVE and the F1 measure (harmonic mean of precision and recall). Generally, MVE can be improved by not making predictions for hard cases, and the F1 penalty is aimed to compensate for this.
On MPI-INF-3DHP and MuPoTS-3D, PCK stands for the percentage of correct keypoints, namely those predicted joints that are within 150 mm of the ground truth.
@@ Appendix C Training Details
Definition of the canonical volume
The canonical space is defined in reference to a specific T-like pose of the default SMPL mesh. All other keypoints (as e.g. shown in the referenced Fig. 4 column 1) are represented in this same coordinate system, as points within this canonical human volume. No explicit conversion between formats is necessary.
@@
The skeleton-based datasets do not provide their correspondence to our SMPL-compatible canonical space. Hence, we perform an approximate initialization and then treat the canonical position of each of several hundred joints (across the skeleton types) as trainable variables during the main training.
One could obtain this initialization in various ways. For this work, we first trained a model for predicting the separate skeleton formats (similar to the separate-heads baseline in [93]). We then ran inference with this predictor on the SURREAL dataset and trained linear regressors to interpolate from SURREAL GT vertices to the predicted keypoints and applied this regressor to the canonical template to obtain the approximate initialization.
The corresponding left and right joints are forced to be symmetrically placed in canonical space. To avoid divergence, we impose a constraint that the canonical positions cannot move away more than 70 mm from their initial positions.
@@
Appendix D Inference Speed Optimization
NLF’s inference-time overhead (for predicting weights through the field MLP) can be eliminated by precomputing the weights once for a chosen set of canonical points. (Typically one wants to localize the same points, i.e. same skeleton formats, for many images.) For reference, in case of NLF-S with no image batching, and about 8000 points to be predicted (mesh vertices and skeletons), forwarding the field MLP to obtain convolution weights takes 7.7 ms, while the rest of the network including the backbone takes 12.7 ms. For NLF-L with batch size 64 the latter takes 587 ms, making the MLP cost negligible in comparison even if we do not precompute it.
Appendix E Uncertainty Estimation
@@
where is the ?stop gradient? operation that blocks the gradient flow during backpropagation. Intuitively, for , multiplying by ?cancels out? the division of by , resulting in the same gradients for as in the case of vanilla Euclidean loss (removing the above-mentioned downweighting of challenging examples), without impacting the gradients for (due to ).
@@
To illustrate that the obtained uncertainties are visually plausible, we show an example in Fig. 5, where the occluded hand has high predicted uncertainty, in line with intuition.
Appendix F Baseline Architecture Comparison without Localizer Field
To emphasize the key importance of using a localizer field in tying together the heatmap-producing convolutional weights for nearby points of the body, we perform the following ablation. We create a baseline architecture which uses no localizer field, and instead defines separate learnable parameters for each SMPL vertex and each joint of each skeleton format.
As seen in Fig. 6, when using the baseline model for inference, the different skeleton format predictions are visibly inconsistent with each other and with the SMPL mesh (see e.g. the H36M arm outside the SMPL body in the lower example), since the weights to localize each point have no enforced relation to each other. This also results in more scattered vertex predictions (see e.g. the hand region). NLF, by contrast, ensures that the different skeletons are localized consistently with each other, and the spatial smoothness of the prediction is improved.
Appendix G Efficient Body Model Fitting
| Iterations | 25 | 37 | 50 | 100 |
| Time | 3 min 18 s | 6 min 54 s | 16 min 35 s | 33 min |
| Error | 14.0 mm | 8.0 mm | 6.2 mm | 5.0 mm |
@@ Appendix G Efficient Body Model Fitting
To further increase efficiency, we may use only a subset of the vertices of the body model during fitting. In Fig. 8, we can see that e.g., by fitting to only 1024 vertices, we increase throughput by a factor of 3, while mIoU only decreases from 0.845 to 0.843. The timing was measured on an NVIDIA RTX 3090 GPU with a batch size of 2048.
In the model transfer experiment described in the main paper, where we fit SMPL-X parameters to clean SMPL mesh vertices with our fitting algorithm, we do not regularize the betas (as the input is a clean mesh without noise), use one iteration and 4096 vertices for fitting (out of the 10475). We also provide a more detailed comparison to the official model transfer code in Tab. 10, showing that our algorithm is much faster even at the same accuracy levels.
@@
We make all code publicly available (main page at https://istvansarandi.com/nlf, repository at https://github.com/isarandi/nlf), including the fitting algorithm (https://github.com/isarandi/smplfitter), to ensure its precise reproducibility.
Appendix H Neural Field Architecture and Details
@@
Appendix I Additional 3DPW Protocol
@@
| Method | MPJPE | P-MPJPE | MVE | P-MVE |
| MeTRAbs-ACAE-S [93] | 60.6 | 41.7 | – | – |
| NLF-S | 59.2 | 41.0 | 71.7 | 52.0 |
| NLF-S +fit | 59.4 | 41.5 | 71.1 | 50.8 |
| MeTRAbs-ACAE-L [93] | 58.9 | 39.5 | – | – |
| NLF-L | 59.2 | 39.5 | 71.0 | 49.2 |
| NLF-L +fit | 59.0 | 40.0 | 70.1 | 48.3 |
Appendix J Further Qualitative Results
In Figs. 10 and 11 we show further results of NLF-L with SMPL-fitting in post-processing (rendered using the Blendify framework [32]).